OpenStudy (anonymous):

What is 5 plus one half?

OpenStudy (anonymous):

5*1/2=5/2 (or 2.5)

OpenStudy (zarkon):

\[5+\frac{1}{2}=\frac{2}{2}*5+\frac{1}{2}=\frac{10}{2}+\frac{1}{2}=\frac{10+1}{2}=\frac{11}{2}=5.5\]

OpenStudy (anonymous):

Zarkon has right! I though that is was multiplication.